OpenStudy (mrhirohito):

Chemical energy is a form of A. neither potential nor kinetic energy. B. potential energy only. C. kinetic energy only. D. both potential and kinetic energy.

OpenStudy (mrhirohito):

i think b T-T

OpenStudy (anonymous):

At point 1: the kinetic energy = 1/2mv12 = 1/2(3kg)(2m/s)2 =6 kg*m2/s2 = 6 N*m = 6 J where m = mass in kilograms v = velocity in m/s m = unit of length, meter s = unit of time, second N = unit of force, Newton 1 N = 1 kg*m/s2 J = unit of energy, Joule 1 J = 1 N*m Also, the potential energy = mgh1 = 0, since h = 0 g = acceleration due to gravity = 9.8m/s2 h1 = height relative to datum (origin), in meters At point 1 the total energy = KE1 + PE1 = 6 + 0 = 6 Joules Note at point 2: PE2 also = 0, therefore KE2 = 6 PE2 = 0 because h2 = 0 since floor is level. If KE2 = KE1 and mass does not change, v2 = v1 = 2m/s Another example Now let's say h1≠h2, let's say h2 = -2m

OpenStudy (anonymous):

h2 = -2m means that point 2 is 2 meters below point 1. A, B, and C represent three possible paths. Note: the path taken has no influence on the solution. Recall that KE = 6 J, PE = 0 J Since total energy remains constant KE1 + PE1 = KE2 + PE2 6 + 0 = KE2 + PE2 We can find PE2 since h2 = -2m PE2 = mgh2 = (3kg)(9.8m/s)(-2m) = -58.8J Now 6 + 0 = KE2 - 58.8 Now solve for KE2 6 + 58.8 = KE2 KE2 = 64.8 J To find v2 KE2 = 1/2mv22 = 1/2(3kg)(v2)2 = 64.8 v22 = 2/3(64.8) = 43.2 v2 = 6.57 m/s An increase of 4.57 m/s

OpenStudy (anonymous):

Finally W1-2 = [7(cos(35°))N](20m) = 114.68J The new equation KE1 + PE1 + W1-2 = KE2 + PE2 6 + 0 + 114.68 = KE2 + 0 120.68 J = KE2 Lastly, to find v2 KE2 = 1/2mv22 120.68 = 1/2(3kg)(v2)2 80.45 = v22 v2 = 8.97 m/s An increase of 6.97 m/s due to applied force across 20 m from point 1 to point 2.

OpenStudy (anonymous):

so B. correct

OpenStudy (mrhirohito):

One more :3

OpenStudy (mrhirohito):

Which of the following statements is true? A. Chemical reactions can neither absorb heat energy nor release heat energy. B. Chemical reactions can only absorb heat energy. C. Chemical reactions can only release heat energy. D. Chemical reactions can either absorb heat energy or release heat energy.

OpenStudy (mrhirohito):

i think D.

OpenStudy (anonymous):

no A.

OpenStudy (mrhirohito):

Thanks

OpenStudy (anonymous):

youre welcome

OpenStudy (anonymous):

Hannah does an experiment. She has two colorless chemicals that she mixes together in a beaker. Which of the following indicates that a chemical change occurs when she mixes the two chemicals together? I. The chemicals change color from colorless to purple. II. A yellow gas forms. III. A brown solid forms in the bottom of the beaker. IV. The temperature of the chemicals rises 15°C.